please help me solve this | x + 6 | = 5
Answer by College Student(505)

The absolute value of an expression always yield a positive result. Therefore, in this equation, the absolute value of 5 and -5 will result in positive 5.

Your answers are: x=-1 and x=-11

Proof: If we substitute these two values in the original equation, we get...

| x + 6 | = 5
| -1 + 6 | = 5
| 5 | = 5

| x + 6 | = 5
| -11 + 6 | = 5
| -5 | = 5